OpenStudy (anonymous):

Subtract the complex numbers. (11 - 7i) - (16 - 11i)

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow 11 - 7i - 16 + 11i }\)

OpenStudy (anonymous):

should i solve from there or is that the answer '

Parth (parthkohli):

More to solve from there.

Parth (parthkohli):

Combine the like terms. That's it.

OpenStudy (anonymous):

11-16=-5 -7i+11i=4i ??

Parth (parthkohli):

What do you get in the answer?

Parth (parthkohli):

Yes

OpenStudy (anonymous):

so = -5+4i

OpenStudy (anonymous):

thanks alot :)

Parth (parthkohli):

\(\huge \color{Purple}{ \mathbb{<tips hat>}}\)